Circular No. 8244

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2001 RG_{100} (LINEAR) CCD exposures taken by A. E. Gleason with the 0.9-m f/3 Spacewatch

reflector at Kitt Peak on Nov. 26.4 UT show that the object 2001 RG_{100} (LINEAR discovery observation from MPS 36374 given below) is definitely a comet, showing a nuclear condensation of diameter ≈ 6′′ and a tail 18′′ long in p.a. 265^{o}. Following a request from the Minor Planet Center, J. Young reports that CCD images taken at Table Mountain on Nov. 27 show a 4′′ coma with a broad, faint tail extending ≈ 12′′ in p.a. 265^{o}.

2001 UT  α_{2000}  δ_{2000}  Mag.
Sept.12.26331  22^{h}45^{m}31.^{s}47  −12^{o}12′40′′.2  19.1

The following orbital elements by B. G. Marsden are from observations published in the MPSs.

Epoch = 2002 July 25.0 TT

T = 2002 July 25.7649 TT  ω = 235.^{o}5880
e = 0.028930  Ω = 137.3131 }2000.0
q = 4.595048 AU  i = 7.9031

a = 4.731943 AU n^{o} = 0.0957513 P = 10.293 years

(65803) 1996 GT P. Pravec, Ondřejov Observatory; L. A. M. Benner, Jet Propulsion

Laboratory (JPL); M. C. Nolan, National Astronomy and Ionosphere Center; P. Kušnirák, Ondřejov Observatory; D. Pray, Greene, RI, U.S.A.; J. D. Giorgini, R. F. Jurgens, and S. J. Ostro, JPL; J.-L. Margot, University of California, Los Angeles; C. Magri, University of Maine at Farmington; A. Grauer, University of Arkansas at Little Rock; and S. Larson, University of Arizona, report that photometric observations — obtained during Nov. 20–24 at Ondřejov, at Steward Observatory, and by Pray — show that (65803) is a binary system with an orbital period of 11.9 hr. The primary rotates with a period of 2.26 hr, and its lightcurve amplitude of 0.10 mag is indicative of a nearly spheroidal shape. Mutual eclipse/occultation events 0.05–0.08 mag deep indicate a secondary-to-primary mean-diameter ratio of 0.2. Arecibo radar delay-Doppler images (2380 MHz, 12.6 cm) obtained on Nov. 23 and 24 independently confirm that (65803) is a binary system. Preliminary estimates of the diameters, based on images at 15-m resolution, are ≈ 800 and ≈ 150 m.

2003 November 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green